OpenStudy (anonymous):

4,Given the function

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[f:x \rightarrow \frac{ 2x }{ x-1 }\],\[x \neq1\] b)state the value of x such that f^-1 does not exist.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

well,first find f inverse

OpenStudy (anonymous):

u meant this?yea?

OpenStudy (anonymous):

the question said that f^-1 is not exist @PFEH.1999

OpenStudy (jhannybean):

The format of your question does not make sense. \(\color{blue}{\text{Originally Posted by}}\) @MARC_ \[f:x \rightarrow \frac{ 2x }{ x-1 }\],\[x \neq1\] b)state the value of x such that f^-1 does not exist. \(\color{blue}{\text{End of Quote}}\) That gets a little confusing to understand :\

OpenStudy (anonymous):

yes @Jhannybean , i can't catch what is the question too ;)

OpenStudy (anonymous):

the answer in the book said it's 2

OpenStudy (anonymous):

but i keep getting =2/0

OpenStudy (jhannybean):

Is it \(f(x) = \dfrac{2x}{x-1}\) ?

OpenStudy (anonymous):

yes ! now i got what is the question (maybe) \[f ^{-1} (x)= \frac{ x }{ x-2 }\]

OpenStudy (anonymous):

so for x=2 does not exist? u meant this?

OpenStudy (anonymous):

it's \[f:x \rightarrow \frac{ 2x }{ x-1 }\]

OpenStudy (jhannybean):

what does the \(:\) and \(\rightarrow\) imply?

OpenStudy (anonymous):

\[f(x)=\frac{ 2x }{ x-1 }\] is the same as \[f:x \rightarrow \frac{ 2x }{ x-1 }\] @Jhannybean

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

maybe start by finding the inverse of f(x)

ganeshie8 (ganeshie8):

\[\large y = \dfrac{2x}{x-1}\] remember how to find the inverse ?

ganeshie8 (ganeshie8):

you find it in two steps : 1) swap x and y 2) solve y

OpenStudy (anonymous):

i got \[f^-1(x)=\frac{x }{ x-2}\]

OpenStudy (anonymous):

correct ;)

ganeshie8 (ganeshie8):

Yes! look at the denominator, what are the excluded values of \(f^{-1}(x)\)

ganeshie8 (ganeshie8):

what value of \(x\) makes the denominator 0 ?

OpenStudy (anonymous):

x=2

ganeshie8 (ganeshie8):

yes why is it a bad value ?

ganeshie8 (ganeshie8):

\(2\) is a bad input for \(f^{-1}(x)\) because it makes the denominator \(0\) and the function goes crazy when denominator approaches 0

ganeshie8 (ganeshie8):

maybe plugin a close number like x = 1.999 and see what the function outputs

ganeshie8 (ganeshie8):

\[\large f^{-1}(x)=\frac{x }{ x-2}\] plugin x=1.999 \[\large f^{-1}(1.999)=\frac{1.999 }{ 1.999-2} =\frac{1.999 }{-0.001} = -1999 \]

ganeshie8 (ganeshie8):

plugin x = 1.99999 \[\large f^{-1}(1.99999)=\frac{1.99999 }{ 1.99999-2} =\frac{1.999 }{-0.00001} = -199999 \]

ganeshie8 (ganeshie8):

as you can see the closer you get to 2, the more crazy huge negative value the function is taking so we say x = 2 is an exclude value for the function

ganeshie8 (ganeshie8):

graph the function in your favorite desmos and see whats going on at x = 2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

in calculas we say when x approaches 2 then f^-1 (x) approaches to \( \infty\)! see the graph:

OpenStudy (anonymous):

* using https://www.desmos.com/calculator

OpenStudy (anonymous):

it's a straight line

OpenStudy (anonymous):

which mean it's undefined right?

OpenStudy (anonymous):

yes , at x=2 then f^-1 (x) is undefined.

OpenStudy (anonymous):

\(\bf\huge\color{#ff0000}{T}\color{#ff2000}{h}\color{#ff4000}{a}\color{#ff5f00}{n}\color{#ff7f00}{k}~\color{#ffaa00}{Y}\color{#ffd400}{o}\color{#bfff00}{u}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) @ganeshie8 @PFEH.1999 @Jhannybean